hello and here is my question... which units would you use to measure the area of a rectangle with 1=1m and w=34cm
Answer by ptaylor(2198)

Well, first, both the length and width must be in either meters or centimeters. We cannot multiply meters and centimeters together and get a useful result. We know that 1 meter is 100 cm and 1 cm is 1/100 meter. We also know that A=L*W, soooo:
We'll work with meters first
A=(1 meter)*(34/100 meter)=34/100 square meters (Remember meter*meter=square meter or meter^2)
Now let's deal in cm
A=(100 cm)*(34 cm)=3400 square centimeters
Now the question is:
Does 34/100 sq meters=3400 square centimeters??
Well, 1 square meter=(100 cm)*(100 cm)=10,000 square cm
So 34/100 square meters=(34/100)*(10,000) square cm and that equals 3400 square centimeters.
So you could use either meters or centimeters and the answer, of course, would be in square meters or square centimeters
Hope this helps---ptaylor